OpenStudy (anonymous):

An equation is shown below: 5x - y = 8 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent.

OpenStudy (anonymous):

@campbell_st I need your help again :D

OpenStudy (campbell_st):

I'd graph the line so show all the solutions... the line is made up of all the points that follow the rule

OpenStudy (campbell_st):

so to get 3 different solutions substitute x -0, x = 1 and x = 2 and find the corresponding y values for an inconsistent solution, you could use a multiply of the line or a parallel line

OpenStudy (anonymous):

@campbell_st I'm confused can you show me how?

OpenStudy (campbell_st):

ok... so if you rewrite the equation you get y = 5x - 8 this is a line you can graph on the number plane let x = 0 y = 5 *0 - 8 so y = -8 the point (0, -8) is on the line x = 1 y = 5*1 - 8 so y = -3 the point (1, -3) is on the line you can do the same for x = 2 to find a point so for the inconsistent line you might choose a line that is parallel or a line that is a multiple 5x - y = 8 multiply every term by 2 what do you get.... then go and graph it

OpenStudy (anonymous):

ok thanks